Find the margin of error from the given 95% confidence interval for the population proportion of patients who said “Drug X was effective in delivering pain relief.” 95% CI: (0.84, 0.96)

Group of answer choices

.03

.06

.12

.90
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.84+  0.96)/ 2=0.9

Margin of error = E = Upper confidence interval - = 0.96 - 0.9=0.06